OpenStudy (anonymous):

One rectangular-shaped piece of metal sliding needs to be cut to cover the exterior of a pole barn. The area of the piece is 30 ft^2. The length is 1 less than 3 times the width. How wide should the metal piece be? Round to the nearest hundredth of a foot. A. 3.33 ft B. 4.3 ft C. 1 ft D. 30 ft

OpenStudy (radar):

Let w = width then 3w -1 = length (given) Area =30 sq ft (given) Area = length times width (3w-1)*(w) = 30 given. Clear the parenthesis 3w^2 -w = 30 Subtract 30 from both sides. 3w^2 -w -30 = 0 Factor the quadratic (3w - 10)(w + 3) = 0 Solve using algebra w + 3 = 0 , then w = -3 Discard as dimension can't be negative 3w - 10 = 0 w=10/3 or 3 1/3 or 3.33 length equal 10 - 1 = 9